# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ G.3 Deep architectures
In App. E.1, we discussed how to derive $p_{f_{0}}\left({x}\right)$ for a neural network with a single hidden layer. In this section, we will discuss the extension of the computation to a network with an arbitrary number of hidden layers. After that, we will focus on the impact that depth has on IGB. In particular, in App. G.3.3 we will show that network depth does not cause IGB, but it amplifies it when the network’s depth increases.
@@ G.3.1 Multi-layer perceptron
The strategy that we will follow is similar to the one presented for the single-hidden-layer counterpart (see App. E.1). We will discuss here the case of MLP with ReLU activation function to focus on the effects of network depth. Extending the results discussed in the single-layer case to deeper networks, including variations such as the inclusion of Max-Pooling, is straightforward. We will propagate the signal across the network layers, keeping track of the changes in the distributions $p_{h^{(l+1)}_{i}}^{(\chi)}\left({x}\right)$ and $p_{\left\langle{h^{(l+1)}_{i}}\right\rangle}\left({x}\right)$. Using the CLT and its extensions considerably simplifies this propagation process because to keep track of changes in the distribution it is sufficient to see how few quantities vary. [^15] This is because a Gaussian distribution is completely determined by the first two cumulants. Compared to the analysis in App. E.1, however, there is an important complicating element. While the elements in the set $\{h^{(1)}_{i}\}$ follow the same distribution (regardless of the choice of the activation function), from the second layer onward the activation function can induce a breaking in symmetry among the layer nodes, in the sense that they will follow, fixed a given configuration for the network weights, different distributions. This symmetry breaking, as we shall see, can cause an accentuation of IGB. To show this point, we will consider in our analysis the ReLU activation function. [^16] We do not include the max pooling to isolate the effect of depth from that of max pooling. Note that absence of max pooling is equivalent of a max pooling with minimum kernel size. Starting from the same setting described in App. E.1, we will derive a set of iterative equations to propagate across multiple layers.
@@
    $\begin{split}&h^{(1)}_{i}\equiv\sum_{j}w^{(0)}_{ij}\xi_{j}\xRightarrow{\text{CLT}}p_{h^{(1)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{d\rightarrow\infty}\mathcal{N}(0,\sigma_{w}^{2})\xRightarrow{\text{ReLU}}p_{g^{(1)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{d\rightarrow\infty}\Theta\left(x\right)\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)+\frac{1}{2}\delta\left(x\right)\,.\end{split}$ (164)
    All the nodes $\{h^{(1)}_{i}\}$ follow the same distribution, and so do the nodes $\{g^{(1)}_{i}\}$ . In particular, this means that
    $\lim_{D\to\infty}\left\langle{g^{(1)}_{i}}\right\rangle=\left\langle{g^{(1)}}\right\rangle,\;\forall i\,.$ (165)
@@
    $h^{(2)}_{i}\equiv\sum_{j}w^{(1)}_{ij}g^{(1)}_{j}\,$. (166)
    It is easy to prove that the generic r.v. $\left(w^{(1)}_{ij}g^{(1)}_{j}\right)$ involved in the sum satisfies the conditions of Eq. (29) and Eq. (30). So, in the limit $N_{1}\rightarrow\infty$, we again have a convergence to a normal distribution. In particular,
    $\begin{split}&p_{h^{(2)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{N_{1}\rightarrow\infty}\mathcal{N}\left(x;\left\langle{g^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}},\textrm{Var}_{\chi}\left({g^{(1)}}\right)S_{\boldsymbol{w}_{i}^{2}}\right)=\mathcal{N}\left(x;\left\langle{g^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}},\textrm{Var}_{\chi}\left({g^{(1)}}\right)\mkern 1.5mu\overline{\mkern-1.5muS_{\boldsymbol{w}_{i}^{2}}\mkern-1.5mu}\mkern 1.5mu\right)\,.\end{split}$ (167)
    In the last step of Eq. (167), as done for $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ in Sec. E.1, we used the concentration result derived in App. C.2.1. In particular that the distribution of the r.v. $S_{\boldsymbol{w}_{i}}\equiv\sum_{j=1}^{N_{1}}w^{(1)}_{ij}$ stays stable in the limit $N_{1}\rightarrow\infty$, [^17] Fluctuations and mean stay both $\mathcal{O}\left(1\right)$. while the distribution of $S_{\boldsymbol{w}_{i}^{2}}\equiv\sum_{j=1}^{N_{1}}\left(w^{(1)}_{ij}\right)^{2}$ asymptotically narrows around the mean value. Note that, as the distribution of $S_{\boldsymbol{w}_{i}}$ does not asymptotically concentrate, and since $\left\langle{g^{(1)}}\right\rangle\neq 0$, each node in the set $\{h^{(2)}_{i}\}$ will follow a different distribution. In particular we will have a set of normally distributed r.v.s, centred on different random points, $\{\left\langle{g^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}}\}_{i}$. After passing through the ReLU activation function, we will have
    $\begin{split}&p_{g^{(2)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{N_{1}\rightarrow\infty}\Theta\left(x\right)\mathcal{N}\left(\left\langle{g^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}},\textrm{Var}_{\chi}\left({g^{(1)}}\right)\mkern 1.5mu\overline{\mkern-1.5muS_{\boldsymbol{w}_{i}^{2}}\mkern-1.5mu}\mkern 1.5mu\right)+\Bigg(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\Bigg(\frac{\left\langle{g^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}}}{\sqrt{2\textrm{Var}_{\chi}\left({g^{(1)}}\right)\mkern 1.5mu\overline{\mkern-1.5muS_{\boldsymbol{w}_{i}^{2}}\mkern-1.5mu}\mkern 1.5mu}}\Bigg)\Bigg)\delta\left(x\right)\,.\end{split}$ (168)
@@
    $p_{h^{(3)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{N_{2}\rightarrow\infty}\mathcal{N}\left(x;\sum_{j=1}^{N_{2}}w^{(2)}_{ij}\left\langle{g^{(2)}_{j}}\right\rangle,\sum_{j=1}^{N_{2}}\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)\right)\,$. (170)
    Note that in this case we cannot take out from the sums $\left\langle{g^{(2)}_{j}}\right\rangle$ and $\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)$ since the nodes $\{g^{(2)}_{j}\}$ are not identically distributed. In App. C.2 we analyze the differences in $p_{h^{(3)}_{i}}^{(\chi)}\left({x}\right)$ between the different nodes $\{h^{(3)}_{i}\}$. In particular, we show that
    $p_{h^{(3)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{N_{2}\rightarrow\infty}\mathcal{N}\left(x;\left\langle{h^{(3)}_{i}}\right\rangle,\textrm{Var}_{\chi}\left({h^{(3)}_{i}}\right)\right)\,$, (171)
    $p_{\left\langle{h^{(3)}_{i}}\right\rangle}\left({x}\right)\xrightarrow{N_{2}\rightarrow\infty}\mathcal{N}\left(x;0,\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(2)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)\,$, (172)
    $p_{\textrm{Var}_{\chi}\left({h^{(3)}_{i}}\right)}\left({x}\right)\xrightarrow{N_{2}\rightarrow\infty}\delta\left(x-\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu\right)\,$. (173)
@@
    $\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({g^{(l-1)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu=\int_{\mathbb{R}}\textrm{Var}_{\chi}\left({g^{(l-1)}_{j}}\right)(x)\;\;\mathcal{N}\left(x;0,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l-2)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)dx\,$, (174)
    $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l-1)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\int_{\mathbb{R}}\left\langle{g^{(l-1)}_{j}}\right\rangle^{2}(x)\;\;\mathcal{N}\left(x;0,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l-2)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)dx.$ (175)
@@
    $p_{h^{(l)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{N_{l-1}\rightarrow\infty}\mathcal{N}\left(x;\left\langle{h^{(l)}_{i}}\right\rangle,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({g^{(l-1)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu\right)\,$, (176)
    $p_{\left\langle{h^{(l)}_{i}}\right\rangle}\left({x}\right)\xrightarrow{N_{l-1}\rightarrow\infty}\mathcal{N}\left(x;0,\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l-1)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)\,.$ (177)
@@
    $\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({g^{(l)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu=\int_{\mathbb{R}}\textrm{Var}_{\chi}\left({g^{(l)}_{j}}\right)(x)\;\;p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)dx\,$, (178)
    $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\int_{\mathbb{R}}\left\langle{g^{(l)}_{j}}\right\rangle^{2}(x)\;\;p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)dx\,.$ (179)
@@
    $\begin{cases}&p_{O^{(c)}}^{(\chi)}\left({x}\right)\xrightarrow{N_{L}\rightarrow\infty}\mathcal{N}\left(x;\left\langle{O^{(c)}}\right\rangle,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({g^{(L)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu\right)\\ &p_{\mu_{c}}^{(\chi)}\left({x}\right)\xrightarrow{N_{L}\rightarrow\infty}\mathcal{N}\left(x;0,\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(L)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)\end{cases}\,.$ (180)
    These two distributions are the only ingredient that we need to replicate the steps described at the end of App. E.1 to get $\mathbb{P}^{\mathcal{W}}\left(f_{0}\right)$.
Figure G.1: Effects induced by the depth of the network. Simulations on an MLP with $L$ hidden layers and ReLU activation function (without any pooling layers) are compared for two different depth values. The theoretical curves are superimposed on the empirical distributions, demonstrating a good agreement between the predictions of our analysis and empirical observations. For a single hidden layer network $(L=1)$ we have a stable distribution (i.e. that does not narrow around $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\mkern-1.5mu}\mkern 1.5mu=0.5$ for $D\rightarrow\infty$) but still peaked on $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\mkern-1.5mu}\mkern 1.5mu=0.5$. Increasing the number of hidden layers the probability mass moves from the center to the extremes of the support. For these simulations we used the GB dataset on model MHLP (see App. I.4 for more details).

### v5

@@ G.3 Deep architectures
In App. E.1, we discussed how to derive $f_{G_{0}}\left({g_{0}}\right)$ for a neural network with a single hidden layer. In this section, we will discuss the extension of the computation to a network with an arbitrary number of hidden layers. After that, we will focus on the impact that depth has on IGB. In particular, in App. G.3.3 we will show that network depth does not cause IGB, but it amplifies it when the network’s depth increases.
@@ G.3.1 Multi-layer perceptron
The strategy that we will follow is similar to the one presented for the single-hidden-layer counterpart (see App. E.1). We will discuss here the case of MLP with ReLU activation function to focus on the effects of network depth. Extending the results discussed in the single-layer case to deeper networks, including variations such as the inclusion of Max-Pooling, is straightforward. We will propagate the signal across the network layers, keeping track of the changes in the distributions $f_{Z^{(l+1)}_{i}}^{(\chi)}\left({z{}{}}\right)$ and $f_{\left\langle{Z^{(l+1)}_{i}}\right\rangle}\left({m}\right)$. Using the CLT and its extensions considerably simplifies this propagation process because to keep track of changes in the distribution it is sufficient to see how few quantities vary. [^15] This is because a Gaussian distribution is completely determined by the first two cumulants. Compared to the analysis in App. E.1, however, there is an important complicating element. While the elements in the set $\{Z^{(1)}_{i}\}$ follow the same distribution (regardless of the choice of the activation function), from the second layer onward the activation function can induce a breaking in symmetry among the layer nodes, in the sense that they will follow, fixed a given configuration for the network weights, different distributions. This symmetry breaking, as we shall see, can cause an accentuation of IGB. To show this point, we will consider in our analysis the ReLU activation function. [^16] We do not include the max pooling to isolate the effect of depth from that of max pooling. Note that absence of max pooling is equivalent of a max pooling with minimum kernel size. Starting from the same setting described in App. E.1, we will derive a set of iterative equations to propagate across multiple layers.
@@
    $\begin{split}&Z^{(1)}_{i}\equiv\sum_{j}w^{(0)}_{ij}\xi_{j}\xRightarrow{\text{CLT}}f_{Z^{(1)}_{i}}^{(\chi)}\left({z{}{}}\right)\xrightarrow{d\rightarrow\infty}\mathcal{N}(0,\sigma_{w}^{2})\xRightarrow{\text{ReLU}}f_{H^{(1)}_{i}}^{(\chi)}\left({h{}{}}\right)\xrightarrow{d\rightarrow\infty}\Theta\left(h{}{}\right)\mathcal{N}\left(h{}{};0,\sigma_{w}^{2}\right)+\frac{1}{2}\delta\left(h{}{}\right)\,.\end{split}$ (164)
    All the nodes $\{Z^{(1)}_{i}\}$ follow the same distribution, and so do the nodes $\{H^{(1)}_{i}\}$ . In particular, this means that
    $\lim_{D\to\infty}\left\langle{H^{(1)}_{i}}\right\rangle=\left\langle{H^{(1)}}\right\rangle,\;\forall i\,.$ (165)
@@
    $Z^{(2)}_{i}\equiv\sum_{j}w^{(1)}_{ij}H^{(1)}_{j}\,$. (166)
    It is easy to prove that the generic r.v. $\left(w^{(1)}_{ij}H^{(1)}_{j}\right)$ involved in the sum satisfies the conditions of Eq. (29) and Eq. (30). So, in the limit $N_{1}\rightarrow\infty$, we again have a convergence to a normal distribution. In particular,
    $\begin{split}&f_{Z^{(2)}_{i}}^{(\chi)}\left({z{}{}}\right)\xrightarrow{N_{1}\rightarrow\infty}\mathcal{N}\left(z{}{};\left\langle{H^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}},\textrm{Var}_{\chi}\left({H^{(1)}}\right)S_{\boldsymbol{w}_{i}^{2}}\right)=\mathcal{N}\left(z{}{};\left\langle{H^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}},\textrm{Var}_{\chi}\left({H^{(1)}}\right)\mkern 1.5mu\overline{\mkern-1.5muS_{\boldsymbol{w}_{i}^{2}}\mkern-1.5mu}\mkern 1.5mu\right)\,.\end{split}$ (167)
    In the last step of Eq. (167), as done for $f_{O_{c}}^{(\chi)}\left({o}\right)$ in Sec. E.1, we used the concentration result derived in App. C.2.1. In particular that the distribution of the r.v. $S_{\boldsymbol{w}_{i}}\equiv\sum_{j=1}^{N_{1}}w^{(1)}_{ij}$ stays stable in the limit $N_{1}\rightarrow\infty$, [^17] Fluctuations and mean stay both $\mathcal{O}\left(1\right)$. while the distribution of $S_{\boldsymbol{w}_{i}^{2}}\equiv\sum_{j=1}^{N_{1}}\left(w^{(1)}_{ij}\right)^{2}$ asymptotically narrows around the mean value. Note that, as the distribution of $S_{\boldsymbol{w}_{i}}$ does not asymptotically concentrate, and since $\left\langle{H^{(1)}}\right\rangle\neq 0$, each node in the set $\{Z^{(2)}_{i}\}$ will follow a different distribution. In particular we will have a set of normally distributed r.v.s, centred on different random points, $\{\left\langle{H^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}}\}_{i}$. After passing through the ReLU activation function, we will have
    $\begin{split}&f_{H^{(2)}_{i}}^{(\chi)}\left({h{}{}}\right)\xrightarrow{N_{1}\rightarrow\infty}\Theta\left(h{}{}\right)\mathcal{N}\left(h{}{};\left\langle{H^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}},\textrm{Var}_{\chi}\left({H^{(1)}}\right)\mkern 1.5mu\overline{\mkern-1.5muS_{\boldsymbol{w}_{i}^{2}}\mkern-1.5mu}\mkern 1.5mu\right)+\Bigg(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\Bigg(\frac{\left\langle{H^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}}}{\sqrt{2\textrm{Var}_{\chi}\left({H^{(1)}}\right)\mkern 1.5mu\overline{\mkern-1.5muS_{\boldsymbol{w}_{i}^{2}}\mkern-1.5mu}\mkern 1.5mu}}\Bigg)\Bigg)\delta\left(h{}{}\right)\,.\end{split}$ (168)
@@
    $f_{Z^{(3)}_{i}}^{(\chi)}\left({z{}{}}\right)\xrightarrow{N_{2}\rightarrow\infty}\mathcal{N}\left(z{}{};\sum_{j=1}^{N_{2}}w^{(2)}_{ij}\left\langle{H^{(2)}_{j}}\right\rangle,\sum_{j=1}^{N_{2}}\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)\right)\,$. (170)
    Note that in this case we cannot take out from the sums $\left\langle{H^{(2)}_{j}}\right\rangle$ and $\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)$ since the nodes $\{H^{(2)}_{j}\}$ are not identically distributed. In App. C.2 we analyze the differences in $f_{Z^{(3)}_{i}}^{(\chi)}\left({z{}{}}\right)$ between the different nodes $\{Z^{(3)}_{i}\}$. In particular, we show that
    $f_{Z^{(3)}_{i}}^{(\chi)}\left({z{}{}}\right)\xrightarrow{N_{2}\rightarrow\infty}\mathcal{N}\left(z{}{};\left\langle{Z^{(3)}_{i}}\right\rangle,\textrm{Var}_{\chi}\left({Z^{(3)}_{i}}\right)\right)\,$, (171)
    $f_{\left\langle{Z^{(3)}_{i}}\right\rangle}\left({m_{z{}{}}}\right)\xrightarrow{N_{2}\rightarrow\infty}\mathcal{N}\left(m_{z{}{}};0,\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(2)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)\,$, (172)
    $f_{\textrm{Var}_{\chi}\left({Z^{(3)}_{i}}\right)}\left({v}\right)\xrightarrow{N_{2}\rightarrow\infty}\delta\left(v-\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu\right)\,$. (173)
@@
    $\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({H^{(l-1)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu=\int_{\mathbb{R}}\textrm{Var}_{\chi}\left({H^{(l-1)}_{j}}\right)(m_{z{}{}})\;\;\mathcal{N}\left(m_{z{}{}};0,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l-2)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)dm_{z{}{}}\,$, (174)
    $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l-1)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\int_{\mathbb{R}}\left\langle{H^{(l-1)}_{j}}\right\rangle^{2}(m_{z{}{}})\;\;\mathcal{N}\left(m_{z{}{}};0,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l-2)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)dm_{z{}{}}.$ (175)
@@
    $f_{Z^{(l)}_{i}}^{(\chi)}\left({z{}{}}\right)\xrightarrow{N_{l-1}\rightarrow\infty}\mathcal{N}\left(z{}{};\left\langle{Z^{(l)}_{i}}\right\rangle,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({H^{(l-1)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu\right)\,$, (176)
    $f_{\left\langle{Z^{(l)}_{i}}\right\rangle}\left({z{}{}}\right)\xrightarrow{N_{l-1}\rightarrow\infty}\mathcal{N}\left(z{}{};0,\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l-1)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)\,.$ (177)
@@
    $\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({H^{(l)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu=\int_{\mathbb{R}}\textrm{Var}_{\chi}\left({H^{(l)}_{j}}\right)(m_{z{}{}})\;\;f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z{}{}}}\right)dm_{z{}{}}\,$, (178)
    $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\int_{\mathbb{R}}\left\langle{H^{(l)}_{j}}\right\rangle^{2}(m_{z{}{}})\;\;f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z{}{}}}\right)dm_{z{}{}}\,.$ (179)
@@
    $\begin{cases}&f_{O_{c}}^{(\chi)}\left({o}\right)\xrightarrow{N_{L}\rightarrow\infty}\mathcal{N}\left(o;\left\langle{O_{c}}\right\rangle,\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({H^{(L)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu\right)\\ &f_{\mu_{c}}^{(\chi)}\left({m}\right)\xrightarrow{N_{L}\rightarrow\infty}\mathcal{N}\left(m;0,\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(L)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\right)\end{cases}\,.$ (180)
    These two distributions are the only ingredient that we need to replicate the steps described at the end of App. E.1 to get $f_{G_{0}}\left({g_{0}}\right)$.
Figure G.1: Effects induced by the depth of the network. Simulations on an MLP with $L$ hidden layers and ReLU activation function (without any pooling layers) are compared for two different depth values. The theoretical curves are superimposed on the empirical distributions, demonstrating a good agreement between the predictions of our analysis and empirical observations. For a single hidden layer network $(L=1)$ we have a stable distribution (i.e. that does not narrow around $\mkern 1.5mu\overline{\mkern-1.5muG_{0}\mkern-1.5mu}\mkern 1.5mu=0.5$ for $D\rightarrow\infty$) but still peaked on $\mkern 1.5mu\overline{\mkern-1.5muG_{0}\mkern-1.5mu}\mkern 1.5mu=0.5$. Increasing the number of hidden layers the probability mass moves from the center to the extremes of the support. For these simulations we used the GB dataset on model MHLP (see App. I.4 for more details).